# Learning Fairer Representations with FairVIC

### v2

@@ 7 Conclusion and Future Work
In this paper, we introduced FairVIC, an in-processing bias mitigation technique that introduces three new terms into the loss function of a neural network- variance, invariance, and covariance. Across our experimental evaluation, FairVIC significantly improves scores for all fairness metrics, with minimal drop in accuracy, compared to previous comparable methods which typically aim to improve only upon a single metric. This balance showcases FairVIC’s strength in providing a robust and effective solution applicable across various tasks and datasets. Future work would look to extend FairVIC to consider multiple protected characteristics simultaneously and expand its utility to image datasets.
8 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Experiment Details
A.1 Dataset Metadata
@@
Detailed metadata for each dataset, including our selection of privileged group, can be found in Table 3. Note that for the language datasets, the number of features is obtained by combining the protected characteristic and the toxicity label with the 50 tokenised text features. For BiasBios, we take architect, attorney, dentist, physician, professor, software engineer, surgeon as the favourable professions, and interior designer, journalist, model, nurse, poet, teacher, and yoga teacher as the unfavourable professions for our binary classification task.
Table 3: Metadata on all four experimental datasets.
@@
Dataset Adult Income COMPAS German Credit CivilComments-WILDS BiasBios
Data modality Tabular Tabular Tabular Text Text
No. of Features 11 8 20 52 52
No. of Rows 48,842 5,278 1,000 50,000 50,000
@@
Target Variable income two_year_recid credit toxicity profession
Favourable Label >50K (1) False (0) Good (1) Non-Toxic (0) Favourable (1)
Unfavourable Label <=50K (0) True (1) Bad (0) Toxic (1) Unfavourable (0)
Protected Characteristic sex race age race gender
Privileged Group male (1) Caucasian (1) >25 (1) white (1) Male (0)
Unprivileged Group female (0) African-American (0) <=25 (0) non-white (0) Female (1)
A.2 Neural Network Configuration
@@
The configurations for the neural networks utilised for both the tabular and language data can be seen in Table 4. To obtain results, each model was run 10 times over random seeds, with a randomised train/test split each time. The averages and standard deviations were then outputted from across all 10 of the runs.
@@
Parameter Tabular Datasets Language Datasets
Neural Network Architecture Dense(128, 64, 32, 2, 32, 64, 128) BiLSTM(64,32), Dense(64, 2, 64)
No. of Epochs 200 50
Batch Size 256 256
Optimiser Adam AdamW
Learning Rate 5e-2 5e-5
Dropout Rate 0.25 0.50
Regularisation L1(1e-4)L2(1e-3) L1(1e-4)L2(1e-3)
@@ Appendix B Full Training Results
In addition to the results and analysis presented in Section 6, this section provides supplementary experiments and figures.
@@ B.1 Feature Importances
In contrast, the FairVIC models for all three datasets demonstrate a strong reduction in the mean importance of protected attributes and proxy variables. This reduction is due to the three additional terms used in FairVIC- variance, invariance, and covariance. We can see that the covariance term exactly minimises the model’s dependency on the protected characteristic, which, in combination with results in Section 6.1, suggests a fairer decision-making process. The reduction in proxy variables should also be noted. Not only does FairVIC successfully reduce the reliance on the protected attribute, but it can also reduce the reliance on any features strongly correlated to the protected attribute. For example, in the Adult Income dataset, sex and relationship have a strong negative correlation ($-0.58$) meaning a model cannot only propagate bias through the use of sex but also through the use of relationship which we see the baseline model rely upon. The FairVIC model sees the mean feature importance of relationship drop by approximately a third and the importance of sex drop by half. This shows FairVIC’s ability to mitigate both direct and indirect biases, leading to more equitable outcomes.
(a) Adult Income dataset.
(b) COMPAS dataset.
(c) German Credit dataset.
Figure 3: Mean feature importances for the baseline and FairVIC models across three tabular datasets. The protected attribute (green) and strong proxy variables to the protected attribute (black) are annotated for their exact feature importance.
@@ B.2 Individual Fairness Results
Following the analysis found in Section 6.2, Table 5 shows the individual fairness on both the baseline and FairVIC with our recommended lambdas, and FairVIC Invariance ($\lambda_{\text{acc}}=0.1,\lambda_{\text{inv}}=0.9,\lambda_{\text{var, cov}}=0.0$) models using their absolute differences to their counterfactual model results. In the Adult Income dataset, the mean absolute difference across all six metrics combined for the baseline model is $\approx 0.0094$, while for FairVIC invariance it is $\approx 0.0055$. In the COMPAS dataset, the mean absolute difference for the baseline model is $\approx 0.0285$, while for FairVIC Invariance it is $\approx 0.0050$. Finally, for the German Credit dataset the mean absolute difference for the baseline model is $\approx 0.0277$, while for FairVIC Invariance it is $\approx 0.0108$. FairVIC’s invariance term, designed to enhance individual fairness, proves to be effective. The FairVIC invariance model consistently achieves significantly absolute differences, demonstrating the success of the approach. In our selection of FairVIC terms, we prioritize group fairness by weighting invariance lower, yet the model still maintains low counterfactual absolute differences.
@@ B.3 Hyperparameter Recommendations
The weights for the loss terms in FairVIC ($\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}}$) were chosen based on insights from our ablation studies.
@@
For the COMPAS and German Credit datasets, the weights were set to $\lambda_{\text{acc, var, inv}}=0.1$ and $\lambda_{\text{cov}}=0.7$. The decision to use a relatively low weight for accuracy ($\lambda_{\text{acc}}=0.1$) stems from the equal ablation study results, which demonstrated that this value achieves the best fairness-accuracy trade-off for these datasets. Group fairness is given significant emphasis, as shown by the higher weight assigned to the covariance term ($\lambda_{\text{cov}}=0.7$), which plays a key role in minimizing disparities across protected groups. Meanwhile, the variance ($\lambda_{\text{var}}$) and invariance ($\lambda_{\text{inv}}$) terms were assigned a weight of 0.1, as this value still allowed for their individual fairness aims to be achieved effectively, thus balancing all fairness and accuracy objectives. Therefore, our default lambda recommendations would be $\lambda_{\text{acc}}=0.1$, $\lambda_{\text{var}}=0.1$, $\lambda_{\text{inv}}=0.1$, and $\lambda_{\text{cov}}=0.7$. These weights were also found to be effective for the CivilComments-Wilds dataset, and thus were utilised here as well.
For the Adult Income dataset, the weights were set to $\lambda_{\text{acc}}=0.2$, $\lambda_{\text{var}}=0.1$, $\lambda_{\text{inv}}=0.1$, and $\lambda_{\text{cov}}=0.6$. A slightly higher weight for accuracy ($\lambda_{\text{acc}}=0.2$) was chosen compared to the COMPAS and German Credit datasets. This decision reflects the findings in Table 6, where a lower accuracy weight ($\lambda_{\text{acc}}=0.1$) led to fairness metrics such as disparate impact exceeding one. While this reflects FairVIC’s ability to actively address fairness concerns, the chosen weights create a careful balance between fairness and predictive performance. FairVIC leverages the increased accuracy weight to account for the complexity of the Adult dataset, characterised by its larger size, while maintaining strong fairness outcomes. The dominant weight assigned to the covariance term ($\lambda_{\text{cov}}=0.6$) further ensures FairVIC prioritizes equitable outcomes across protected groups, achieving robust group fairness without compromising individual fairness or accuracy. The BiasBios dataset also require stronger interventions from the covariance term to improve group fairness, therefore we assign weights of $\lambda_{\text{acc, var, inv}}=0.05$, and $\lambda_{\text{cov}}=0.85$ to this dataset.
These weight configurations reflect the flexibility of FairVIC in balancing dataset-specific requirements for individual fairness, group fairness, and prediction accuracy. To effectively utilise FairVIC, we recommend users prioritise their specific fairness objectives—whether group fairness or individual fairness—based on the desired application context. If group fairness is the primary goal, assigning a higher weight to the covariance term ($\lambda_{\text{cov}}$) can help mitigate disparities across protected groups. Conversely, for tasks requiring equitable treatment at the individual level, increasing the weight of the invariance ($\lambda_{\text{inv}}$) term will enhance individual fairness. For users aiming to optimize specific performance metrics, such as accuracy or fairness, we could recommend conducting a grid search over the loss term weights ($\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}}$) to fine-tune the trade-offs for their dataset.
Additionally, users should consider the complexity and size of their dataset when configuring the weights. Larger datasets or those with greater feature diversity may require higher weights for accuracy ($\lambda_{\text{acc}}$) to maintain strong predictive performance. Fairness requirements may also vary depending on the level of societal or organisational impact, and we encourage users to carefully assess the implications of their choices in real-world deployments. Finally, we recommend evaluating FairVIC’s performance using a range of fairness and accuracy metrics to ensure that the selected configuration aligns with the intended goals of the application.
@@ B.4 Model Representation Analysis
An example latent space visualisation from the baseline model and FairVIC can be seen in Figure 4. In the baseline model, we observe a separation between subgroups, where women (subgroup 0) are predominantly located in the upper region and men (subgroup 1) in the lower region of the latent space. This separation suggests that the baseline model’s representations may be influenced by the protected attribute, leading to the biased decision-making reported in Table 1. In contrast, the FairVIC model shows a more condensed and overlapping distribution of both subgroups within the same latent space. This indicates, alongside results in Table 1 and feature importance in Figure 3(a) that FairVIC has successfully reduced the model’s reliance on the protected characteristic and any proxy variables, thereby promoting more equitable representations. The overlapping and compact structure in the FairVIC latent space demonstrates that similar data points, regardless of their subgroup membership, are mapped closer together, ensuring that the model’s predictions are not unfairly biased towards one group over the other.
@@ B.5 Model Optimization Analysis
B.6 Theoretical Analysis
In this section, we theoretically analyze FairVIC and show how each individual loss term is sub-differentiable.
@@ B.6.1 Theorem 1
$L_{\text{var}}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\gamma-\sigma(z))$ (4)
where $z$ is the latent embeddings for the input $x$. $z=g_{\theta}(x)$ is continuous in $\theta$, where $g_{\theta}$ is the function/layer that maps input $x$ to the latent embedding $z$. $\sigma(z)$ is the standard deviation of a continuous variable in a finite sample, which is continuous except at rare instances where all $z_{j}$ are identical. Even in this degenerate case, $\sigma(\cdot)$ is sub-differentiable. The $\text{max}(0,\cdot)$ operator is only non-differentiable at 00, where the sub-derivative set is $[0,1]$. Hence $\text{max}(0,\cdot)$ is sub-differentiable w.r.t $\theta$. The invariance term is defined as:
@@
$L_{\text{cov}}=\frac{\sqrt{\sum_{i=1}^{N}\left(\left(\hat{y}_{i}-\mathbb{E}[\hat{y}]\right)^{\top}\cdot P_{i}\right)^{2}}}{N}$ (6)
where $\sum_{i=1}^{N}\left(\left(\hat{y}-\mathbb{E}[\hat{y}]\right)^{\top}\cdot P\right)^{2}$ is a sum of squares, which is smooth and differentiable. The square root is differentiable for non-zero input and sub-differentiable at 00.

### v3

@@ 7 Conclusion and Future Work
In this paper, we introduced FairVIC, an in-processing bias mitigation technique that introduces three new terms into the loss function of a neural network- variance, invariance, and covariance. Across our experimental evaluation, FairVIC significantly improves scores for all fairness metrics, with minimal drop in accuracy, compared to previous comparable methods which typically aim to improve only upon a single metric. This balance showcases FairVIC’s strength in providing a robust and effective solution applicable across various tasks and datasets. Future work will extend FairVIC to multiple protected attributes and targets.
Declaration on Generative AI
The author(s) have not employed any Generative AI tools.
@@ Appendix A Experiment Details
A.1 FairVIC Training Algorithm
During training, the model iterates over epochs $E$, with the data shuffled into batches. For each batch, the model produces predictions $\hat{Y}$, which are compared with the true labels $Y$ using a suitable accuracy loss function (e.g., binary cross-entropy, hinge loss, or Huber loss). The resulting loss is then minimised by an optimiser.
The pseudocode below summarises the full training procedure for FairVIC. It integrates the accuracy loss alongside the proposed fairness-promoting objectives- variance, invariance, and covariance- as defined in Section 4.1. The total loss is computed as a weighted combination of these terms. Subsequently, gradients are computed, and the optimiser adjusts the model parameters with respect to this combined loss.
The multipliers $\lambda$ enable users to balance the trade-off between fairness and predictive performance, which is typical in bias mitigation techniques. Assigning a higher weight to $\lambda_{\text{acc}}$ directs the model to prioritise accuracy, while increasing the weights of $(\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}})$ shifts the focus towards enhancing fairness in the model’s predictions. In our implementation, the lambda coefficients $(\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}})$ are constrained such that their sum equals one. In other words, $\lambda_{\text{acc}}=1-\lambda_{\text{var}}-\lambda_{\text{inv}}-\lambda_{\text{cov}}$. This normalisation ensures the optimisation will not produce multiple solutions in the form $\{k\cdot\lambda_{\text{acc}},k\cdot\lambda_{\text{var}},k\cdot\lambda_{\text{inv}},k\cdot\lambda_{\text{cov}}\},\;k\in\mathbb{R}$.
Algorithm 1 FairVIC Loss Function
1Input: Model $M$, Epochs $E$, Batch size $B$, Data $D$, Protected attribute $P$, Weights ($\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}}$)
2Output: Trained Model $M$
3Initialise $M$
4for $e\in\{1,\dots,E\}$ do
5  Shuffle data $D$
6  for each batch $\{(X,Y)\}\in D$ with size $B$ do
7   $\hat{Y}\leftarrow M(X)$
8   $Z\leftarrow\text{BottleneckLayer}(X)$
9   Calculate FairVIC Loss:
[…]
11    $L_{\text{var}}\leftarrow\text{VarianceLoss}(Z)$
12    $L_{\text{inv}}\leftarrow\text{InvarianceLoss}(\hat{Y},M(\text{Flip}(X,P)))$
13    $L_{\text{cov}}\leftarrow\text{CovarianceLoss}(\hat{Y},P)$
14    $L_{\text{total}}\leftarrow\lambda_{\text{acc}}L_{\text{acc}}+\lambda_{\text{var}}L_{\text{var}}+\lambda_{\text{inv}}L_{\text{inv}}+\lambda_{\text{cov}}L_{\text{cov}}$
15   Compute gradients $\nabla L_{\text{total}}\leftarrow\frac{\partial L_{\text{total}}}{\partial M}$
16   Update model parameters $M\leftarrow M-\alpha\nabla L_{\text{total}}$
A.2 Dataset Metadata
@@
Detailed metadata for each dataset, including our selection of privileged group, can be found in Table 3. Note that for the language datasets, the number of features is obtained by combining the protected characteristic and the target label with the 50 tokenised text features and for image datasets, this is the pixels in an image plus the protected attribute and target label. For BiasBios, we take architect, attorney, dentist, physician, professor, software engineer, surgeon as the favourable professions, and interior designer, journalist, model, nurse, poet, teacher, and yoga teacher as the unfavourable professions for our binary classification task.
Table 3: Metadata on all seven experimental datasets.
@@
Dataset Adult Income COMPAS German Credit CivilComments-WILDS BiasBios CelebA UTKFace
Data Modality Tabular Tabular Tabular Text Text Image Image
No. of Features 11 8 20 52 52 128x128 +2 128x128 +2
No. of Rows 48,842 5,278 1,000 50,000 50,000 50,000 23,705
@@
Target Variable income two_year_recid credit toxicity profession Blond_\_Hair age
Favourable Label >50K (1) False (0) Good (1) Non-Toxic (0) Favourable (1) Not Blond (-1) <30 (1)
Unfavourable Label <=50K (0) True (1) Bad (0) Toxic (1) Unfavourable (0) Blond (1) >=30 (0)
Protected Characteristic sex race age race gender Male race
Privileged Group male (1) Caucasian (1) >25 (1) white (1) Male (0) Male (1) White (1)
Unprivileged Group female (0) African-American (0) <=25 (0) non-white (0) Female (1) Female (0) Non-White (0)
A.3 Neural Network Configuration
@@
The configurations for the neural networks utilised for the tabular, language, and image data can be seen in Table 4. To obtain results, each model was run 10 times over random seeds, with a randomised train/test split each time. The averages and standard deviations were then outputted from across all 10 of the runs.
@@
Parameter Tabular Datasets Language Datasets Image Datasets
Network Architecture Dense(128, 64, 32, 2, 32, 64, 128) BiLSTM(64,32) + Dense(64, 2, 64) Conv2D(32, 64)+ Dense(128, 64, 32, 2, 32, 64, 128)
No. of Epochs 200 50 50
Batch Size 256 256 256
Optimiser Adam AdamW AdamW
Learning Rate 5e-2 5e-5 5e-5
Dropout Rate 0.25 0.50 0.25
Regularisation L1(1e-4)L2(1e-3) L1(1e-4)L2(1e-3) L1(1e-4)L2(1e-3)
@@ Appendix B Full Training Results
In addition to the results and analysis presented in Section 6, this section provides supplementary experiments and figures. First, the qualitative visualisations for the COMPAS and German Credit can be seen in Figure 3, following the analysis of the Adult Income dataset in Figure 1. Discussion on these results can be seen in Section 6.1.
(a) COMPAS dataset.
(b) German Credit dataset.
Figure 3: A qualitative analysis of the absolute differences from the ideal value (e.g., perfect accuracy and fairness) in performance (left) and fairness (right) metrics of comparable techniques on the COMPAS and German Credit datasets.
@@ B.1 Feature Importances
In contrast, the FairVIC models for all three datasets demonstrate a strong reduction in the mean importance of protected attributes and proxy variables. This reduction is due to the three additional terms used in FairVIC- variance, invariance, and covariance. We can see that the covariance term exactly minimises the model’s dependency on the protected characteristic, which, in combination with results in Section 6.1, suggests a fairer decision-making process. The reduction in proxy variables should also be noted. Not only does FairVIC successfully reduce the reliance on the protected attribute, but it can also reduce the reliance on any features strongly correlated to the protected attribute. For example, in the Adult Income dataset, sex and relationship have a strong negative correlation (0.58-0.58) meaning a model cannot only propagate bias through the use of sex but also through the use of relationship which we see the baseline model rely upon. The FairVIC model sees the mean feature importance of relationship drop by approximately a third and the importance of sex drop by half. This shows FairVIC’s ability to mitigate both direct and indirect biases, leading to more equitable outcomes. On the COMPAS dataset, while race remains as the second most important feature, its actual importance dropped by $\approx 41\%$.
Figure 4: Mean feature importances derived from SHAP values for the baseline and FairVIC models across the three tabular datasets. The protected attribute (green) and its strong proxy variables (black) are annotated with their exact feature importance values.
@@ B.2 Individual Fairness Results
Following the analysis found in Section 6.2, Table 5 shows the individual fairness on both the baseline and FairVIC with our recommended lambdas, and FairVIC Invariance ($\lambda_{\text{acc}}=0.1,\lambda_{\text{inv}}=0.9,\lambda_{\text{var, cov}}=0.0$) models using their absolute differences to their counterfactual model results. In the Adult Income dataset, the mean absolute difference across all six metrics combined for the baseline model is $0.0094$, while for FairVIC invariance it is $0.0055$. In the COMPAS dataset, the mean absolute difference for the baseline model is $0.0285$, while for FairVIC Invariance it is $0.0050$. Finally, for the German Credit dataset the mean absolute difference for the baseline model is $0.0277$, while for FairVIC Invariance it is $0.0108$. FairVIC’s invariance term, designed to enhance individual fairness, proves to be effective. The FairVIC invariance model consistently achieves significantly absolute differences, demonstrating the success of the approach. In our selection of FairVIC terms, we prioritize group fairness by weighting invariance lower, yet the model still maintains low counterfactual absolute differences.
@@ B.3 Hyperparameter Recommendations
The weights for the loss terms in FairVIC ($\lambda_{\text{acc}},\lambda_{\text{var}},\lambda_{\text{inv}},\lambda_{\text{cov}}$) were chosen based on insights from our ablation studies. We have outlined the weights we used in our evaluation in Table 6.
Table 6: Our final lambda selection and recommendations for each of the seven datasets evaluated upon.
Adult Income Compas German Credit CivilComments- WILDS BiasBios CelebA UTKFace
$\lambda_{\text{acc}}$ 0.20 0.10 0.10 0.10 0.05 0.10 0.20
$\lambda_{\text{var}}$ 0.10 0.10 0.10 0.10 0.05 0.05 0.10
$\lambda_{\text{inv}}$ 0.10 0.10 0.10 0.10 0.05 0.05 0.10
$\lambda_{\text{cov}}$ 0.60 0.70 0.70 0.70 0.85 0.80 0.60
To help users configure FairVIC effectively, we provide a recommended starting point and guidance for adapting the loss term weights to suit different fairness and performance goals. While our paper focuses on generalisable defaults, in a real-world deployment, one could envisage a tuning process to identify optimal results for a given application or domain.
@@
We recommend starting with the configuration $\lambda_{\text{acc}}=0.1$, $\lambda_{\text{var}}=0.1$, $\lambda_{\text{inv}}=0.1$, and $\lambda_{\text{cov}}=0.7$, which provided strong results across multiple datasets including COMPAS, German Credit, and CivilComments-WILDS. This balanced setting encourages individual fairness, supports diverse representations, and strongly targets group fairness without significantly compromising accuracy. The decision to use a relatively low weight for accuracy ($\lambda_{\text{acc}}=0.1$) stems from the equal ablation study results, which demonstrated that this value achieves the best fairness-accuracy trade-off for these datasets. Group fairness is given significant emphasis, as shown by the higher weight assigned to the covariance term ($\lambda_{\text{cov}}=0.7$), which plays a key role in minimizing disparities across protected groups. Meanwhile, the variance ($\lambda_{\text{var}}$) and invariance ($\lambda_{\text{inv}}$) terms were assigned a weight of 0.1, as this value still allowed for their individual fairness aims to be achieved effectively, thus balancing all fairness and accuracy objectives.
To adjust for different dataset characteristics or application goals, users should consider the intended role of each loss term:
If accuracy needs improvement while fairness is already strong, increasing the accuracy term weight can help the model prioritise predictive performance. For instance, in the Adult Income and UTKFace datasets, increasing $\lambda_{\text{acc}}$ to 0.2 (while slightly lowering $\lambda_{\text{cov}}$) led to better trade-offs. This is likely because both datasets have relatively high-quality features and well-separated class distributions, allowing the model to benefit from a greater emphasis on discriminative capacity once basic fairness has been addressed.
When group fairness is a higher priority, the covariance term should be weighted more heavily. This term explicitly penalises statistical dependence between predictions and group membership, and is most impactful when datasets show strong baseline disparities between protected and privileged groups. For example, image-based datasets such as CelebA often contain visually encoded group cues that strongly correlate with target labels, while text-based datasets like BiasBios may exhibit linguistic patterns linked to social or demographic attributes. In such cases, increasing $\lambda_{\text{cov}}$ (e.g. to 0.8 or 0.85) helps reduce spurious correlations and promotes more group-independent predictions.
To enhance individual fairness, a higher weight on the invariance term helps the model maintain consistency under counterfactual changes to the protected attribute. This is especially useful for datasets where protected features are not strongly entangled with the label, such that counterfactual consistency is a plausible fairness goal. Although our default already includes invariance, users with specific fairness requirements may find that increasing $\lambda_{\text{inv}}$ offers more robust guarantees.
In practice, the ideal configuration depends on the dataset size, complexity, and the fairness-performance trade-off required by the application. Larger or more complex datasets may require a slightly higher weight on the accuracy term to avoid underfitting, while fairness-sensitive domains may justify higher weights on the fairness terms.
  • Default configuration: $\lambda_{\text{acc}}=0.1,\ \lambda_{\text{var}}=0.1,\ \lambda_{\text{inv}}=0.1,\ \lambda_{\text{cov}}=0.7$
  • To improve accuracy: Increase $\lambda_{\text{acc}}$
  • To emphasise group fairness: Increase $\lambda_{\text{cov}}$
  • To emphasise individual fairness: Increase $\lambda_{\text{inv}}$
If users have a specific optimisation goal—such as maximising a particular fairness or accuracy metric—a targeted grid search over the loss weights may be appropriate, though this falls outside the primary scope of our work.
@@ B.4 Model Representation Analysis
An example latent space visualisation from the baseline model and FairVIC can be seen in Figure 5. In the baseline model, we observe a separation between subgroups, where women (subgroup 0) are predominantly located in the upper region and men (subgroup 1) in the lower region of the latent space. This separation suggests that the baseline model’s representations may be influenced by the protected attribute, leading to the biased decision-making reported in Table 1. In contrast, the FairVIC model shows a more condensed and overlapping distribution of both subgroups within the same latent space. This, alongside results in Table 1 and feature importance in Figure 4, indicates that FairVIC has successfully reduced the model’s reliance on the protected characteristic and any proxy variables, thereby promoting more equitable representations. The overlapping and compact structure in the FairVIC latent space demonstrates that similar data points, regardless of their subgroup membership, are mapped closer together, ensuring that the model’s predictions are not unfairly biased towards one group over the other.
@@ B.5 Model Optimization Analysis
B.6 Theoretical Analysis and Discussion
In this section, we theoretically analyse FairVIC and show how each individual loss term is sub-differentiable.
@@ B.6.1 Theorem 1
$L_{\text{var}}=\frac{1}{d}\sum_{j=1}^{d}\max(0,\gamma-\sigma_{j}(z))$ (5)
where $z$ is the latent embeddings for the input $x$. $z=g_{\theta}(x)$ is continuous in $\theta$, where $g_{\theta}$ is the function/layer that maps input $x$ to the latent embedding $z$. $\sigma(z)$ is the standard deviation of a continuous variable in a finite sample, which is continuous except at rare instances where all $z_{j}$ are identical. Even in this degenerate case, $\sigma(\cdot)$ is sub-differentiable. The $\text{max}(0,\cdot)$ operator is only non-differentiable at 0, where the sub-derivative set is $[0,1]$. Hence $\text{max}(0,\cdot)$ is sub-differentiable w.r.t $\theta$. The invariance term is defined as:
@@
$\color[rgb]{0,0,0}L_{\text{cov}}=\frac{1}{N}\sqrt{\sum_{i=1}^{N}\left(\left(\hat{y}_{i}-\mathbb{E}[\hat{y}]\right)\cdot P_{i}\right)^{2}}$} (7)
where $\sum_{i=1}^{N}\left(\left(\hat{y}-\mathbb{E}[\hat{y}]\right)\cdot P\right)^{2}$ is a sum of squares, which is smooth and differentiable. The square root is differentiable for non-zero input and sub-differentiable at 0.
@@
B.7 Trade-Off Discussion
The design of FairVIC is grounded in the recognition that group fairness metrics such as Statistical Parity Difference, Equalized Odds Difference, and Disparate Impact capture distinct and often conflicting notions of fairness. Foundational theoretical work by Kleinberg et al. [41] has shown that, under realistic conditions- such as differing base rates between groups- it is impossible for any prediction system to simultaneously satisfy multiple fairness criteria, such as calibration and equalized odds, without introducing trade-offs.
FairVIC does not attempt to satisfy these fairness criteria simultaneously in a formal sense. Rather than optimising for any specific fairness definition directly, FairVIC introduces inductive biases at the representation level through variance, invariance, and covariance regularisation. These objectives are agnostic to downstream fairness metrics and aim to induce feature representations that are disentangled from protected attributes while preserving predictive signal.
The improvements observed across multiple fairness metrics are therefore not the result of directly encoding incompatible fairness constraints, but an emergent property of the learned representations. As shown in Table 1, FairVIC achieves significant reductions in Statistical Parity Difference and Equalized Odds Difference across all datasets, and improves Disparate Impact in most cases. In the case of our experiments, these fairness improvements come with modest reductions in predictive performance- consistent with expected trade-offs. Sometimes, improvements to one fairness metric such as Disparate Impact- slightly affect the results for another fairness metric, highlighting that such trade-offs are context-dependent and not inevitable.
Importantly, FairVIC does not commit to a normative fairness definition during training. This design choice reflects a practical and ethical consideration: in many real-world settings, there may be no consensus on which fairness metric best captures the relevant notion of harm or justice. By focusing on representation-level properties rather than metric-specific constraints, FairVIC supports post hoc evaluation across multiple fairness metrics, encouraging empirical pluralism and transparency in fairness assessments.
In summary, FairVIC does not circumvent impossibility theorems, nor does it attempt to. Rather, it sidesteps the need to encode conflicting fairness notions directly, and instead fosters representation learning that generalises well across subgroups. The observed metric improvements - and their trade-offs - are empirical outcomes of this strategy, not theoretical contradictions.